# Moment-based parameter inference with error guarantees for stochastic reaction networks

### v1

@@ III.2 Partial Data Case
A hierarchy of outer approximations analogous to (21), but with partial moment observations, can then be defined as
$\psi^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\mathbb{R}_{+}^{(N_{d}\times R)}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}^{obs}_{l}\leq\bm{z}^{obs}_{j}\leq k_{j}\hat{\bm{y}}^{obs}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ \end{array}}\right\}.$ (27)
@@
$\min_{\psi^{d}}\bm{y^{\bm{\alpha}}}\leq\bm{y}^{\bm{\alpha}*}\leq\max_{\psi^{d}}\bm{y^{\bm{\alpha}}},$ (28)
where $\bm{y}^{\bm{\alpha}}$ denotes one of the unobserved moments and $\bm{y}^{\bm{\alpha}*}$ denotes its true value.
@@ IV Numerical results: Application to Biochemical Reaction Networks
In this section, we apply our approach to biochemical models using synthetic data (complete or partially observed). In the complete data case, we first use the Schlögl model as an introductory example and study the tightness of the parameter bounds in dependence on the moment order and sample size of the data. We show in detail how to form a constrained set and then compare the bounds on rate parameters obtained from intervals of moments with those based on the true moments. Secondly, we consider the genetic toggle switch with complete data again to illustrate that our method can integrate datasets for inferring common parameters. As a final example, we consider synthetic datasets with unobserved species obtained from a model of post-transcriptional gene regulation and demonstrate that datasets with different unobserved species can be combined together to increase the tightness of parameter bounds. We also demonstrate how to bound moments of unobserved species.
In all examples, the assumed prior knowledge of the parameters is that at least one parameter is known, i.e., the matrix $C$ is a diagonal matrix with at least one non-zero term and the corresponding value in $c$ is the true parameter value. The datasets used in this section are obtained by collecting samples at the same time point after the burn-in period of $n$ trajectories of Gillespie’s stochastic simulation algorithm (Gillespie, 1977), and assumed to be at steady state. We then use 2000 bootstrap samples of these data points to compute 95% bootstrap confidence intervals of each moment. We then consider the constrained sets $\zeta^{d}$ (21) and $\psi^{d}$ (27) by setting up the semidefinite programs using the package YALMIP (Lofberg, 2004) and carrying out the optimisation with the solver Mosek (MOSEK ApS, 2022) and default parameters.
IV.1 Schlögl Model with Full Observations
@@
Figure 2: Parameter inference of the Schlögl model with varying moment order and sample size. (a) Optimisation-based bounds on $\langle x^{1}\rangle_{\pi}$ and $\langle x^{2}\rangle_{\pi}$ converge to the exact moment (abbr. to mmt.) values. Methods see (Kuntz et al., 2019). (b) Bootstrap intervals of moments, with sample sizes 10000 and 20000, and the true moment values. Values are log-transformed with base 10. (c) Bounds on $k_{1}$ (left) and $k_{3}$ (right) obtained from $zeta^{d}$ using bootstrap moment intervals with sample sizes 10000 and 20000, and using the exact moments. (d) Feasible regions of $k_{1}$ and $k_{3}$ obtained from moment intervals with sample size 20000 for $d=3,\ldots,7$. (e) Feasible regions of $k_{1}$ and $k_{3}$ obtained from moment intervals with sample sizes varying from 5000 to 25000 and $d=4$. True parameters are $\bm{k}=(2,3,1,4)^{\top}$ and $k_{2}$ and $k_{4}$ are assumed to be known in (b) - (e).
@@
IV.2 Toggle Switch Model with Multiple Full Observations
Next, we demonstrate our approach for parameter inference with rational propensities from multiple perturbation datasets. We consider the toggle switch model Gardner, Cantor, and Collins (2000) with two chemical species and rational propensities:
@@ IV.2 Toggle Switch Model with Multiple Full Observations
IV.2.1 Post-Transcriptional Gene Regulation with Partial Observations
Finally, we apply our method to networks with unobserved species, which we refer to as the partial data case. As a simplified model of post-transcriptional regulation in cells (Statello et al., 2021; Hausser and Zavolan, 2014) with bimolecular reactions, we consider the following 2-species model:
@@ IV.2.1 Post-Transcriptional Gene Regulation with Partial Observations
Figure 4: Parameter bounds of the post-transcriptional regulation model with partial data. (a) Relative bounds on the unknown parameters when $k_{1},k_{3},k_{5}$ are known (left) and when $k_{2},k_{4},k_{5}$ are known (right) with $d=4$. In each case, we compare the bounds obtained when $X_{1}$ is unobserved (abbr. to unobs.); when $X_{2}$ is unobserved; when we have two datasets each has only one species measured; and when we have full data. (b) Histogram of $X_{1}$ and $X_{2}$ showing they have similar magnitudes at steady states. (c) Relative bounds on parameters obtained with $d=3,4,5$. We compare the bounds on the unknown parameters when assuming $k_{3},k_{4},k_{5}$ known but $X_{1}$ observed (abbr. to obs.) and when assuming $k_{1},k_{2},k_{5}$ known but $X_{2}$ observed. In all panels, the true parameters are $\bm{k}=(6,0.8,5,0.5,1)^{\top}$. The sample size is $n=20000$ in the first three panels.
@@
$\min_{\psi^{d}}\bm{y^{\bm{\alpha}}}\leq\bm{y}^{\bm{\alpha}*}\leq\max_{\psi^{d}}\bm{y^{\bm{\alpha}}}$ (39)
where $\bm{y}^{\bm{\alpha}}$ denotes one of the unobserved moments and $\bm{y}^{\bm{\alpha}*}$ denotes its true value. The bounds on these moments are divided by the mean of bootstrap intervals for comparison (Fig. 5). Although the quality of the bounds decreases with the order of the moment, the bounds on the means or cross moments are relatively tight and constrain the moments of the latent species reasonably well. This demonstrates that our approach provides quantified uncertainties for parameters and moment predictions of stochastic reaction networks with latent species.
@@ V Discussion
We proposed a novel inference approach for stochastic reaction networks based on moment constraints and mathematical programming. We used the moment equations and the moment matrices to form a constrained set of the rate parameters and outer approximate this set by introducing variables to replace the non-linear terms in the constraints. Upper and lower bounds on the parameters can then be obtained by optimising over this set, forming a convex semidefinite programming problem with a unique optimum. Our method takes intervals of the stationary moments as input, and the bounds obtained are guaranteed to contain the true parameters under the assumption that these intervals include the true moments.
In contrast to likelihood-based and simulation-based approaches, our approach avoids the solution of the forward problem, i.e., predicting moments and distributions for given parameters, and thus avoids intractable likelihood computations or computationally expensive simulations associated with these approaches. Unlike likelihood- and simulation-based approaches, our approach is guaranteed to converge in polynomial time and its computational cost scales only with the number of parameters and the order of the constrained set (Traub and Woźniakowski, 1982; Vandenberghe and Boyd, 1996). Moreover, our approach enjoys the advantage that it provides bounds on the parameters and thus provides error guarantees under mild assumptions.
@@
Our approach represents a computationally tractable and efficient convex outer approximation to the original optimisation problem, which itself falls into the broad and complicated category of non-convex quadratically constrained quadratic programs. We have demonstrated that our approximation generally provides tight and informative bounds in a number of applications. Yet it remains unclear what conditions guarantee the tightness of our bounds. Another limitation is that our approach is based on raw moments, but other statistics could be advantageous to improve the sufficiency, convergence properties, and computational costs of our parameter inference. Further work will be needed to investigate these issues.
The present method can be extended to include observation error models (see Appendix C). Another possible extension is towards transient dynamics by considering time-dependent moment equations in which the moments must also satisfy a hierarchy of positive semidefinite constraints (Sakurai and Hori, 2018; Dowdy and Barton, 2018b; Holtorf and Barton, 2023). While our approach focuses on parameter inference, there are several other applications. For example, our approach could be used in: (i) identifiability analysis through identifying unbounded parameters; (ii) model selection as rejecting models with unfeasible parameter regions; and (iii) model reduction achieved by pruning reactions whose rate bounds include zeros.
@@
MATLAB code will be made available at github.com/pthomaslab/SDP-inference.

### v2

@@ III.2 Partial Data Case at Steady State
A constrained set analogous to $\zeta^{d}$ in (28), but with partial moment observations, can then be defined as
$\psi^{d}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}\in{\mathbb{R}_{+}^{N_{d}}}\\ Z\in\mathbb{R}_{+}^{N_{d}\times R}\end{array}}{\begin{array}[]{|ll}k_{j}\hat{\bm{y}}^{obs}_{l}\leq\bm{z}^{obs}_{j}\leq k_{j}\hat{\bm{y}}^{obs}_{u},&j=1,\ldots,R\\ \text{trace}(A^{d}_{\bm{\alpha}}Z)=0,&|{\bm{\alpha}}|\leq d-d_{b}+1\\ \bm{z}_{j}^{\top}\bm{h}_{:N_{d}}=k_{j},&j=1,\ldots,R\\ M_{d}^{s}(\bm{y})\succeq 0,&s=0,\ldots,S\\ ZC^{\top}=\bm{y}\bm{c}^{\top}\\ \end{array}}\right\}.$ (34)
@@
$\min_{\psi^{d}}y^{\bm{\alpha}}\leq y^{\bm{\alpha}*}\leq\max_{\psi^{d}}y^{\bm{\alpha}},$ (35)
where $y^{\bm{\alpha}}$ denotes one of the unobserved moments and $y^{\bm{\alpha}*}$ denotes its true value.
@@
III.3 Generalisation to Time-resolved Data
Inference from time-resolved data follows analogously to the steady-state case using the transient moment equations (20), which provide relations between the parameters and the generalised moments. Let $\mathcal{P}$ be the selected set of values of $\rho$ from which we compute intervals $(\hat{\bm{\bar{y}}}_{l}(\rho),\hat{\bm{\bar{y}}}_{u}(\rho))$ of the generalised rational moments in (18) from time-dependent data. Additionally, at the start and end points ($t=0,T$) intervals of the rational moments $(\hat{\bm{y}}_{l}(t),\hat{\bm{y}}_{u}(t))$ are used depending on whether these time points are known or observed.
We create sets of auxiliary variables corresponding to the cross-terms:
$\bar{Z}(\rho)\coloneqq\bm{\bar{y}}(\rho)(k_{0}\ \bm{k}^{\top})=(\bm{\bar{z}}_{0}(\rho)\ \cdots\ \bm{\bar{z}}_{R}(\rho))\in\mathbb{R}_{+}^{N_{d}\times(R+1)},$ (36)
[…]
A suitable outer approximation can then be written as the following SDP:
$\tau^{d}_{\mathcal{P}}=\left\{{\begin{array}[]{l}\bm{k}\in\mathcal{K}\\ \bm{y}(0)\in{\mathbb{R}_{+}^{N_{d}}}\\ \bm{y}(T)\in{\mathbb{R}_{+}^{N_{d}}}\\ \text{for }\rho\in\mathcal{P}:\\ \quad\bar{Z}(\rho)\in\mathbb{R}_{+}^{N_{d}\times(R+1)}\end{array}}{\begin{array}[]{|ll}\text{for }t=0,T:\\ \quad\hat{\bm{y}}_{l}(t)\leq\bm{y}(t)\leq\hat{\bm{y}}_{u}(t),&\text{if available}\\ \quad\bm{y}^{\top}(t)\bm{h}_{:N_{d}}=1\\ \quad M_{d}^{s}(\bm{y}(t))\succeq 0,&s=0,\ldots,S\\ \text{for }\rho\in\mathcal{P}:\\ \quad k_{j}\hat{\bm{\bar{y}}}_{l}(\rho)\leq\bm{\bar{z}}_{j}(\rho)\leq k_{j}\hat{\bm{\bar{y}}}_{u}(\rho),&j=0,\dots,R\\ \quad{{\bm{\bar{h}}}_{\bm{\alpha}}^{\top}}(\bm{y}(T)-e^{\rho T}\bm{y}(0)+\rho\bm{\bar{z}}_{0}(\rho))=\text{trace}(A^{d}_{\bm{\alpha}}\bar{Z}_{1:R}(\rho)),&|\bm{\alpha}|\leq d-d_{h}-d_{b}+1\\ \quad\bm{\bar{z}}_{j}(\rho)^{\top}\bm{h}_{:N_{d}}=k_{j}\frac{e^{\rho T}-1}{\rho}&j=0,\ldots,R\\ \quad M_{d}^{s}(\bm{\bar{z}}_{0}(\rho))\succeq 0,&s=0,\ldots,S\\ \quad\bar{Z}_{1:R}(\rho)C^{\top}=\bm{\bar{z}}_{0}(\rho)\bm{c}\end{array}}\right\}$ (37)
The first inequality represents rational moment interval data at the start and end points, and the second and third lines represent moment and SDP constraints that they satisfy. The following lines constrain the generalised rational moments on a set of test points that satisfy the transient moment equations. The constraints on $\bm{\bar{z}}_{j}(\rho)^{\top}\bm{h}_{:N_{d}}$ follow from the corresponding constraints $\bm{y}^{\top}(t)\bm{h}_{:N_{d}}=1$ and linearity of integration. Note that, if $\rho=0$, the limit needs to be taken, namely $\bm{\bar{z}}_{j}(0)^{\top}\bm{h}_{:N_{d}}=k_{j}T$. Similar conditions for the generalised moments, including the SDP inequalities, were obtained in Ref. Dowdy and Barton (2018b).
@@ IV Numerical results: Application to Biochemical Reaction Networks
In this section, we apply our approach to biochemical models using synthetic data (complete or partially observed). In the complete data case, we first use the Schlögl model as an introductory example and study the dependence of parameter bounds on the moment order and sample size. We show how to form a constrained set and then compare the bounds on rate parameters obtained from intervals of moments with those based on the true moments. Secondly, we consider the genetic toggle switch with complete data to illustrate that our method can integrate datasets for inferring common parameters. We then consider synthetic data with unobserved species obtained from a model of post-transcriptional gene regulation and demonstrate that datasets with different unobserved species can be combined to tighten parameter bounds and estimate the moments of the unobserved species. Finally, we consider a simple birth-death model with transient data observed. We show that both parameters can be bounded simultaneously with time-course data, and the quality of the bounds depends on the choice of summary statistics and the sample size.
In all steady-state examples, the assumed prior knowledge of the parameters is that at least one parameter is known, i.e., the matrix $C$ is a diagonal matrix with at least one non-zero term and the corresponding value in $c$ is the true parameter value. The datasets used in this section are obtained by collecting samples at the same time point after the burn-in period of $n$ trajectories of Gillespie’s stochastic simulation algorithm (Gillespie, 1977), and assumed to be at steady state. We then use 2000 bootstrap samples of these data points to compute 95% bootstrap confidence intervals of each moment. In the transient setting, data simulation is described explicitly in Sec. IV.4. We then consider the constrained sets by setting up the semidefinite programs using the package YALMIP (Lofberg, 2004) and carrying out the optimisation with the solver Mosek (MOSEK ApS, 2022) and default parameters.
IV.1 Schlögl Model with Full Steady-State Observations
@@
Figure 2: Parameter inference of the Schlögl model with varying moment order and sample size. (a) Optimisation-based bounds on $\langle x^{1}\rangle_{\pi}$ and $\langle x^{2}\rangle_{\pi}$ converge to the exact moment (abbr. to mmt.) values. Methods see (Kuntz et al., 2019). (b) Bootstrap intervals of moments, with sample sizes 10000 and 20000, and the true moment values. Values are log-transformed with base 10. (c) Bounds on $k_{1}$ (left) and $k_{3}$ (right) obtained from $\zeta^{d}$ using bootstrap moment intervals with sample sizes 10000 and 20000, and using the exact moments. (d) Feasible regions of $k_{1}$ and $k_{3}$ obtained from moment intervals with sample size 20000 for $d=3,\ldots,7$. (e) Feasible regions of $k_{1}$ and $k_{3}$ obtained from moment intervals with sample sizes varying from 5000 to 25000 and $d=4$. True parameters are $\bm{k}=(2,3,1,4)^{\top}$ and $k_{2}$ and $k_{4}$ are assumed to be known in (b) - (e).
@@
IV.2 Toggle Switch Model with Multiple Full Observations at Steady State
Next, we demonstrate our approach for the case of parameter inference with multiple perturbation datasets. Perturbations here describe parameter changes due to conditions that represent experimental changes in gene induction levels or CRISPR perturbations in cells. We consider the toggle switch model Gardner, Cantor, and Collins (2000) with two chemical species and rational propensities:
@@ IV.2 Toggle Switch Model with Multiple Full Observations at Steady State
IV.3 Post-Transcriptional Gene Regulation with Partial Observations at Steady State
We apply our method to networks with unobserved species, which we refer to as the partial data case. As a simplified model of post-transcriptional regulation in cells (Statello et al., 2021; Hausser and Zavolan, 2014) with bimolecular reactions, we consider the following 2-species model:
@@ IV.3 Post-Transcriptional Gene Regulation with Partial Observations at Steady State
Figure 4: Parameter bounds of the post-transcriptional regulation model with partial data. (a) Relative bounds on the unknown parameters when $k_{1},k_{3},k_{5}$ are known (left) and when $k_{2},k_{4},k_{5}$ are known (right) with $d=4$. In each case, we compare the bounds obtained when $X_{1}$ is unobserved; when $X_{2}$ is unobserved; when we have two datasets each with only one species measured; and when we have full data. (b) Histogram of $X_{1}$ and $X_{2}$ showing they have similar magnitudes at steady state. (c) Relative bounds on parameters obtained with $d=3,4,5$. We compare the bounds on the unknown parameters when assuming $k_{3},k_{4},k_{5}$ known but $X_{1}$ observed and when assuming $k_{1},k_{2},k_{5}$ known but $X_{2}$ observed. In all panels, the true parameters are $\bm{k}=(6,0.8,5,0.5,1)^{\top}$. The sample size is $n=20000$ in panels (a) and (c).
@@
$\min_{\psi^{d}}y^{\bm{\alpha}}\leq y^{\bm{\alpha}*}\leq\max_{\psi^{d}}y^{\bm{\alpha}}$ (48)
where $y^{\bm{\alpha}}$ denotes one of the unobserved moments and $y^{\bm{\alpha}*}$ denotes its true value. The bounds on these moments are divided by the mean of bootstrap intervals for comparison (Fig. 5). Although the quality of the bounds decreases with the order of the moment, the bounds on the means or cross moments are relatively tight and reasonably constrain the moments of the latent species. This demonstrates that our approach provides quantified uncertainties for parameters and moment predictions of stochastic reaction networks with latent species.
@@
IV.4 Birth-Death Process with Time-dependent Observations
To illustrate the use of our method for transient data, we consider a simple birth-death process:
$a_{2}]{a_{1}}X,$ (49)
with mass-action propensity functions
$a_{1}(x)=k_{1},\ a_{2}(x)=k_{2}x.$ (50)
We generate test data by sampling time courses using the SSA starting from $x=0$. To obtain intervals on the generalised moments, we take observations at sample time points $t\sim\text{Uniform}(0,T)$ (Fig. 6a). We then use bootstrapping to obtain intervals on the generalised moments $\bm{\bar{y}}(\rho)$ for a selected set of values of $\rho$, and on the moments at the endpoints.
We observe that the feasible regions of $k_{1}$ and $k_{2}$ are bounded tightly around the true parameters and show a positive correlation between the two parameters (Fig. 6b). When moment data at the endpoint are used, more constraints are introduced, and the feasible region shrinks compared to generalised moments only. To study the influence of the summary statistics on the tightness of the bounds, we consider different sets of $\rho$ taken as combinations of $\rho=0,1,-1$ (Fig. 6c). We see that the parameters cannot be bounded from above for $\rho=-1$ while they are bounded for non-negative values (e.g. $\rho=0$ or $\rho=1$). Including more values of $\rho$ tightens bounds, and the best bounds are obtained when combining all three values.
In Fig. 6d, we study the convergence of our bounds with sample size. The bounds approach the true parameters as the sample size is increased, and they converge faster as more $\rho$ values are included. To demonstrate this, we use the fact that the transient solution of the CME is a time-varying Poisson distribution and compute the generalised moments analytically (nominally corresponding to an infinite sample size). Only for $\rho=0$, the bounds do not converge to the true parameters even with the exact values of the generalised moments. Intuitively, negative values of $\rho$ put more weight at the end of the time series; positive values emphasise the transients at the beginning; and zero values average across time. Hence a suitable range of $\rho$ needs to be selected on a case-by-case basis considering the time scale of the dynamics. Our findings highlight that the choice of summary statistics influences identifiability, bound tightness, and convergence.
Figure 6: Parameter inference of a birth-death model with time-dependent data. (a) Data simulated from the birth-death model (49) with sampling time uniformly distributed between 0 and 10. These data are used to obtain intervals on the generalised moments. An example trajectory of the dynamic is also plotted. (b) Feasible regions of $k_{1}$ and $k_{2}$ with and without 5000 additional observations at the end-point $T$ using three values of $\rho=0,1,-1$. (c) Bounds on both parameters with different choices of $\rho$. Dashed lines indicate that the parameter is unbounded above. (d) The log relative tightness for different sample sizes, which is calculated as $\text{log}((k_{u}-k_{l})/k+1)$ where $k_{u}$ and $k_{l}$ are the upper and lower bounds we obtain for a parameter $k$. A missing value of the tightness indicates that the parameter is unbounded above. Infinite sample size ($\infty$”) corresponds to inference performed with exact moments computed from the time-varying Poisson solution. In panels (a) - (c), the sample size of the transient data is 10000. In panels (c) and (d), no data at the end-point are considered. The true parameters are $\bm{k}=(5,1)^{\top}$ and $T=10$ in all panels and all optimisation was done with $d=3$ with intervals obtained from $2,000$ bootstrap samples.
@@ V Discussion
We proposed a novel inference approach for stochastic reaction networks based on moment constraints and mathematical programming. We used the moment equations and the moment matrices to form a constrained set of the rate parameters and outer approximate this set by introducing variables to replace the non-linear terms in the constraints. Upper and lower bounds on the parameters can then be obtained by optimising over this set, forming a convex semidefinite programming problem with a unique optimum. Our method takes intervals of the stationary moments or the generalised moments as input, and the bounds obtained are guaranteed to contain the true parameters under the assumption that these intervals include the true moments.
In contrast to likelihood-based and simulation-based approaches, our approach avoids the solution of the forward problem, i.e., predicting moments and distributions for given parameters, and thus avoids intractable likelihood computations or computationally expensive simulations associated with these approaches. When the solution is known, our result converges to the interval of the method of moments estimates with bootstrap resampling, but our approach has the advantage that the moment equations do not need to be solved explicitly. Unlike likelihood- and simulation-based approaches, our approach is guaranteed to converge in polynomial time and its computational cost scales only with the number of parameters and the order of the constrained set (Traub and Woźniakowski, 1982; Vandenberghe and Boyd, 1996). Moreover, our approach enjoys the advantage that it provides bounds on the parameters and thus provides error guarantees under mild assumptions.
@@
Our approach represents a computationally tractable and efficient convex outer approximation to the original optimisation problem, which falls into the broad and complicated category of non-convex quadratically constrained quadratic programs. We have demonstrated that our approximation generally provides tight and informative bounds in a number of applications. Yet it remains unclear what conditions guarantee the tightness of our bounds. For example, it is likely that for inference problems with partial observations, products of unknown parameters and unobserved moments cannot be identified using the linearisation involved in our outer approximations. In such cases, the original non-convex inference problem can produce tighter bounds with a different optimisation solver and potentially much larger computational cost. Furthermore, the present method can be extended to include observation errors by coupling true and noise-corrupted moments (see Appendix C). Unless the hyperparameters of the noise model are known, estimating the additional noise parameters leads to a non-convex cubically constrained program that can also suffer from poor identifiability under linearisation. Further work will be needed to investigate these issues.
An extension of our method would be to utilise moment intervals other than bootstrap intervals as input data. For example, one could compute concentration inequalities to couple tail probabilities with moment equations. Another possible development would be to consider parameter distributions instead of point estimates of parameters. Such parameter heterogeneity arises, for example, due to extrinsic noise in cells Zechner et al. (2012); Dattani and Barahona (2017); Thomas and Shahrezaei (2021) and can be implemented in our framework by adding extra static species, which are not changed in reactions. Inference of moments of these parameters then proceeds by producing bounds on moments of the unobserved species analogue to the partial data case (Fig. 5). The corresponding forward problem that bounds moments given known parameter distributions has been discussed by Sakurai and Hori(Sakurai and Hori, 2022), and the same constraints could be used in our inverse problem to infer bounds on moments of the parameter distributions. This approach introduces additional variables and equations for cross-moments of parameters and species, thus adding to the complexity of the optimisation problem.
Here, we employed raw moments for measurements at steady state and generalised moments for time-course inference. In the latter case, generalised moments serve as summary statistics of time courses where the values of the exponent $\rho$ control the weight given to early and late time points. We have shown that these exponents must be chosen carefully and evaluated case-by-case to obtain tight parameter bounds. We also note that the use of generalised moments differs from instantaneous moments taken at discrete time intervals in standard setups (Zechner et al., 2012; Ruess and Lygeros, 2015; Davidović et al., 2022). Still, the generalised moments can be interpreted as a time-averaged version of the instantaneous ones when these are sufficiently frequently sampled and thus provide similar information. It remains an open question whether employing other statistics could improve the sufficiency, convergence properties and computational cost of our parameter inference method.
Our method could be further used to explore the design space of synthetic biochemical circuits within the context of the work of Sakurai and Hori  Sakurai and Hori (2018a). Our approach could be applied to that problem if one uses the desired output moments of a synthetic circuit as moment intervals. Optimising over parameters then provides bounds on the feasible parameter space, thus circumventing the simulation effort needed for the forward problem when designing circuits with defined statistical properties.
@@
MATLAB code is available at github.com/pthomaslab/SDP-inference.